# GRPose: Learning Graph Relations for Human Image Generation with Pose Priors

### v1

@@ Introduction
Datasets. We evaluated our model on the Human-Art (Ju et al. 2023a) and LAION-Human (Ju et al. 2023c) datasets. The Human-Art dataset comprises 50,000 high-quality images from 5 real-world and 15 virtual scenarios, featuring human bounding boxes, key points and textual descriptions. The LAION-Human dataset consists of approximately 1 million image-caption pairs, filtered by high image quality and human estimation confidence scores, using a diverse range of human activities and more realistic images. On LAION-Human dataset we randomly selected 200,000 samples for training and 20,000 samples for testing.
@@
Implementation Details. For a fair comparison, our diffusion pipeline adopts the same Stable Diffusion 1.5111https://huggingface.co/runwayml/stable-diffusion-v1-5 as previous methods. We used the Adam optimizer (Kingma and Ba 2014) with an initial learning rate of $1\times 10^{-5}$. We trained models on 8×8$\times$8 × NVIDIA L40S-48GB, using the Pytorch and PytorchLightning [^2] https://lightning.ai/docs/pytorch/stable/. The batch size was set as 6, and the number of epochs was 20. The weight of pose loss $\alpha$ was set as 0.01 and we adopted a gradual constraint strategy, adding pose perception loss in the last 5 epochs. We replaced the text prompts with an empty string with a probability of 0.5. For the PGI, we set the parameter $K$ of the KNN algorithm to 9. During inference, the number of steps $T$ for the DDIM sampler was set to 50. We employed mmpose [^3] https://github.com/open-mmlab/mmpose framework as the pretrained pose estimation networks. It is worth noting that, in the evaluation phase we used a different pre-trained pose estimation network during inference process to assess the accuracy of the pose in the generated images. This ensures the independence and reliability of the evaluation.
@@
Metrics. We adopted Average Precision (AP), Similarity Average Precision (SAP) and Person Count Error (PCE) (Cheong, Mustafa, and Gilbert 2022). Higher scores in AP and SAP indicate better pose alignment. Additionally, we employed the Fréchet Inception Distance (FID) (Heusel et al. 2017) and Kernel Inception Distance (KID) (Bińkowski et al. 2018) to assess the quality of the generated images, and used the CLIP-Score (Radford et al. 2021) to evaluate the alignment between the image and text.
@@
Figure 4: Results of Multi-Pose Generation. Our model outperforms ControlNet in generating multiple poses.
Comparison with SOTA Methods
@@
Quantitative results comparing our method with other state-of-the-art (SOTA) approaches are shown in Table 1. The experimental results indicate that our GRPose achieved the highest AP and SAP. On the Human-Art dataset, our model attained AP and SAP of 49.50% and 70.84% respectively, bringing an improvement of 9.98% in AP compared to ControlNet. Similarly, on the LAION-Human dataset, our model showed a consistent improvement of 6.06% in AP compared to HumanSD (Ju et al. 2023b). Notably, our method significantly improved pose alignment accuracy, as evidenced by the AP, SAP, and PCE metrics. We observed that SD1.5 using only a text prompt without pose conditions lacked the ability for pose alignment. Although controllable diffusion models such as ControlNet and T2I-Adapter used pose conditions for spatial control, their capabilities of pose alignment were limited due to incorporating it without considering the intrinsic high-order associations between different pose parts. While retaining pose alignment, our model also maintained good performance in image quality and text alignment with minor decline in scores.
@@
Figure 5: Comparison of different base diffusion models for our GRPose. We compared SD1.5, Anime Art and Realistic models of different styles. GRPose effectively enhances each model with superior pose alignment.
@@
We further compared the visualizations with other methods (Zhang, Rao, and Agrawala 2023; Mou et al. 2024; Ju et al. 2023b; Rombach et al. 2022) across several samples, as shown in Figure 7. We found that SD1.5 lacked pose alignment capability due to rely on solely text prompt. While T2I-Adapter and ControlNet exhibited good text semantic ability, their pose accuracy was poor, and HumanSD also showed differences in pose alignment. In contrast, our GRPose demonstrated superior pose alignment, significantly improving the details in human image generation. We also verified the model’s ability to handle multiple poses in a single condition, as shown in Figure 4. Compared to ControlNet, our model maintained accurate alignment with multiple poses, proving the feasibility of the graph learning approach. Our model addressed issues such as unnatural body positions and misaligned pose, significantly enhancing the accuracy of pose estimation. We performed transfer verification of the base model across different SD 1.5 variants from the open-source community Civitai, as shown in Figure 5. Our GRPose, when applied to different base models, enhanced their pose-guided alignment capabilities, demonstrating that GRPose is a plug-and-play solution for various diffusion models.
@@
Figure 6: Visual comparison of PGI. Our model with PGI demonstrates enhanced pose guidance due to graph associations.
@@
Table 2: Results of each component. $\mathcal{L}_{p}$ represents the Pose Perception Loss. These results verify the effectiveness of each component in our framework.
@@
Upper-Body ControlNet 15.51 66.45 0.63
@@
Full-Body ControlNet 30.78 70.12 1.29
@@
Figure 7: Visual comparison of ours and other methods. The samples are from the Human-Art dataset, with each row representing a sample along with its corresponding pose and prompt. Our GRPose, HumanSD, ControlNet, and T2I-Adapter were provided with both a prompt and a pose, while SD1.5 was provided with only a prompt. Our model achieved outstanding results in visual quality and pose alignment.
Table 4: Results of different graph number. The use of graph convolution layers at different stages significantly improves pose alignment. This experiment represents our empirical exploration.
@@
We conducted experiments on the Human-Art dataset to validate the effectiveness of each component, as shown in Table 2. Compared to ControlNet, the addition of the Pose Perception Loss $\mathcal{L}_{p}$ led to an increase of 4.29% in AP and 1.18% in SAP, demonstrating that the pose perception loss effectively constrained the pose differences in the feature space from an alternative perspective. Incorporating the PGI further improved performance of our model, with AP and SAP increasing by 8.28% and 1.44% respectively. This improvement indicated that PGI, through its integration of complex graph topological relationships, accurately captured high-order relationships between pose parts. It was evident that after PGI associating the information of different pose parts, our model produced significant improvements in both AP and SAP, addressing previous issues such as blurred limbs and unrealistic poses, as shown in Figure 4.
To verify the ability of generating portraits, we conducted additional experiments on upper body and full body samples from the cartoon subset on the Human-Art dataset respectively, as shown in Table 3. In the upper body scenario, our model outperformed ControlNet by 16.36% in the AP, while in the full body scenario it was 8.35% higher in the AP. The results showed that compared to ControlNet, our model had more outstanding performance in upper body generation than in full body generation, further verifying the effectiveness of our PGI associations. The performance in the upper body scenario indicated that our model achieved good results in portrait generation, which demonstrates the high potential in many downstream applications.
@@
We conducted empirical exploration to assess the performance of our model with varying numbers of graph structures applied to the latent, pose, and fused features, as shown in Table 4. The results indicated a significant performance increase as the number of graphs increased from 1 to 4. We did not report more results due to computational resource limitations.
In addition, we visualized the change of pose alignment with and without PGI, as shown in Figure 6. In the first row of samples, the leg pose showed a noticeable error and was partially missing without PGI. In the second row, the arm was in an incorrect pose. With PGI, pose guidance significantly improved image quality, bringing the generated pose of the image closer to the origin. This demonstrates that capturing high-order relationships through the graph structure effectively facilitates information propagation between different pose parts.
@@
We examined the impact of different loss weights $\alpha$ for the Pose Perception Loss $\mathcal{L}_{p}$, as shown in Table 5. It can be found that the optimal loss weight is $0.01$. However, when the weight was set to 0.01, there was a significant decrease in AP, suggesting that it imposed constraints on the diffusion loss. A larger weight for the pose perception loss may cause the model to learn interference in the feature space.
Conclusion
In this paper, we proposed a framework, named Graph Relation Pose (GRPose), for pose-guided human image generation. Unlike to previous methods that inject pose priors into diffusion models based on Euclidean space, we used GCNs to capture the higher ordeer relationships between different pose parts. Following the ControlNet approach, our method utilizes both a text prompt and a pose prior to generate human images. We designed a Progressive Graph Integrator (PGI), which adopts a hierarchical structure within the Adapter to ensure that higher-order associations gradually correct the pose. Additionally, we proposed a novel pose perception loss, which uses a pre-trained pose estimation network to constrain the network, further enhancing the accuracy of pose guidance. Our approach achieves SOTA performance on the Human-Art and LAION-Human datasets with the best performance in the accuracy of pose alignment.

### v2

@@ 4.1 Experimental Settings
Datasets. We evaluated our model on the Human-Art (Ju et al. 2023a) and LAION-Human (Ju et al. 2023c) datasets. The Human-Art dataset comprises 50,000 high-quality images from 5 real-world and 15 virtual scenarios, featuring human bounding boxes, key points and textual descriptions. The LAION-Human dataset consists of approximately 1 million image-caption pairs, filtered by high image quality and human estimation confidence scores, using a diverse range of human activities and more realistic images. In the LAION-Human dataset, we randomly selected 200,000 samples for training and 20,000 samples for testing.
@@
Implementation Details. For a fair comparison, our diffusion pipeline adopts the same Stable Diffusion 1.5111https://huggingface.co/runwayml/stable-diffusion-v1-5 as previous methods. We used the Adam optimizer (Kingma and Ba 2014) with an initial learning rate of $1\times 10^{-5}$. We trained models on 8×8$\times$8 × NVIDIA L40S-48GB, using Pytorch and PytorchLightning [^2] https://lightning.ai/docs/pytorch/stable/. The batch size was set as 6, and the number of epochs was 20. The weight of pose loss $\alpha$ was set to 0.01 and we adopted a gradual constraint strategy, adding pose perception loss in the last 5 epochs. We replaced the text prompts with an empty string with a probability of 0.5. For the PGI, we set the parameter $K$ of the KNN algorithm to 9. During inference, the number of steps $T$ of the DDIM sampler was set to 50. We used the MMpose [^3] https://github.com/open-mmlab/mmpose framework as the pre-trained pose estimation networks. It is worth noting that in the evaluation phase we used a different pre-trained pose estimation network during the inference process to assess the accuracy of the pose in the generated images. This ensures the independence and reliability of the evaluation.
@@
Metrics. We adopted Average Precision (AP), Similarity Average Precision (SAP) and Person Count Error (PCE) (Cheong et al. 2022). Higher scores in AP and SAP indicate better pose alignment. Furthermore, we used the Fréchet Inception Distance (FID) (Heusel et al. 2017) and the Kernel Inception Distance (KID) (Bińkowski et al. 2018) to assess the quality of the generated images and used the CLIP-Score (Radford et al. 2021) to assess the alignment between the image and text.
@@
Figure 4: Two cases of Multi-Pose Generation. Our model outperforms ControlNet in generating multiple poses.
4.2 Comparison with SOTA Methods
@@
Quantitative results comparing our method with other state-of-the-art (SOTA) approaches are shown in Table 1. The experimental results indicate that our GRPose achieved the highest AP and SAP. On the Human-Art dataset, our model attained AP and SAP of 49.50% and 70.84% respectively, bringing an improvement of 9.98% in AP, compared with ControlNet. Similarly, on the LAION-Human dataset, our model showed a consistent improvement of 6.06% in AP compared to HumanSD (Ju et al. 2023b). Notably, our method significantly improved pose alignment accuracy, as evidenced by the AP, SAP, and PCE metrics. We observed that SD1.5 using only a text prompt without pose conditions lacked the ability for pose alignment. Although controllable diffusion models such as ControlNet and T2I-Adapter used pose conditions for spatial control, their capabilities of pose alignment were limited due to incorporating it without considering the intrinsic high-order associations between different pose parts. While retaining pose alignment, our model also maintained good performance in image quality and text alignment with minor decline in scores.
@@
Figure 5: Qualitative Results of our GRPose with different base diffusion models. We compared SD1.5, Anime Art and Realistic models of different styles.
@@
We further compared the visualized qualitative results with other methods (Zhang et al. 2023; Mou et al. 2024; Ju et al. 2023b; Mou et al. 2024; Rombach et al. 2022) across several samples, as shown in Figure 7. We found that SD1.5 lacked effective pose alignment capability since it relies solely on text prompts. Although T2I-Adapter and ControlNet exhibited a good ability to comprehend textual semantics, their performance of pose alignment was poor, and HumanSD also performs poorly in pose alignment. In contrast, our GRPose demonstrated good pose alignment, significantly enhancing the details in human image generation. We also verified the model’s ability in handling multiple poses in a single condition, as shown in Figure 4. Compared with ControlNet, our model can achieve accurate alignment with multiple poses, proving the feasibility of the graph learning approach. Our model addressed issues such as unnatural body positions and misaligned pose, significantly enhancing the accuracy of pose estimation. We perform transfer verification of the base model across different variants of SD-1.5 from the open-source community Civitai, as shown in Figure 5. Our GRPose, when integrated with different base models, can effectively improve the performance of pose-guided alignment, demonstrating that GRPose is a plug-and-play and effective solution.
@@
Figure 6: Effect of PGI. Our model with PGI demonstrates better pose guidance due to graph associations.
@@
Table 2: Results of each component. $\mathcal{L}_{p}$ denotes the Pose Perception Loss. The results verify the effectiveness of each component in our framework.
@@
Upper Body ControlNet 15.51 66.45 0.63
@@
Full Body ControlNet 30.78 70.12 1.29
@@
Figure 7: Qualitative comparison between ours and other methods. The samples are from the Human-Art dataset, where each row illustrates a sample along with its corresponding pose and prompt.
Table 4: Effects of different graph number. The use of graph convolution layers at different stages significantly improves the quality of pose alignment.
@@ 4.3 Ablation Studies
We conducted experiments on the Human-Art dataset to validate the effectiveness of each component, as shown in Table 2. Compared with ControlNet, the usage of the Pose Perception Loss $\mathcal{L}_{p}$ led to an increase of 4.29% in AP and 1.18% in SAP, demonstrating that the pose perception loss effectively constrained the pose differences in the feature space from an alternative perspective. Incorporating the PGI further improved the performance of our model, with AP and SAP increasing by 8.28% and 1.44% respectively. This improvement indicated that PGI, benefited from its modeling of complex graph topological relationships, accurately captured high-order relationships between pose parts. It is evident that after PGI associates the information of different parts of the pose, our model produced significant improvements in both AP and SAP, addressing previous issues such as blurred limbs and unrealistic poses, as shown in Figure 4.
To verify the ability of portrait generation, we conducted additional experiments on upper body and full body samples from the cartoon subset in Human-Art, as shown in Table 3. In the upper body scenario, our model outperformed ControlNet by 16.36% in the AP. In the full body scenario it was 8.41% higher in the AP. The results showed that compared with ControlNet, our model performs much better in upper body generation than in full body generation, further verifying the effectiveness of our PGI associations. The performance in the upper body scenario indicated that our model achieved good results in portrait generation, which demonstrates the high potential in many downstream applications.
@@
We conducted empirical exploration to assess the performance of our model with varying numbers of graph structures, as shown in Table 4. The results indicated that we can observe a significant performance increase when the number of graph layers increased from 1 to 4. We did not report more results due to computational resource limitations. In addition, we visualize our results of pose alignment with and without PGI in Figure 6. In the first row of samples, the leg pose showed a noticeable error and was partially missing without PGI. In the second row, the arm was in an incorrect pose. With PGI, pose guidance significantly improved image quality, bringing the generated pose of the image closer to the origin. This shows that capturing high-order relationships through the graph structure effectively facilitates information propagation between different parts of the pose.
@@
We investigated the impact of different loss weights $\alpha$ on Pose Perception Loss $\mathcal{L}_{p}$, as shown in Table 5. It can be found that the optimal loss weight is $0.01$. However, when the weight was set to 0.1, there was a significant performance decrease in AP, suggesting that it imposed over-high constraints on the diffusion loss. A higher weight for pose perception loss may lead to a failure of pose perception.
5 Conclusion
In this paper, we proposed a framework named Graph Relation Pose (GRPose), for pose-guided human image generation. We designed a Progressive Graph Integrator (PGI), which adopts a hierarchical structure within the Adapter to ensure that higher-order associations gradually correct the pose. Additionally, we proposed a novel pose perception loss, which uses a pre-trained pose estimation network to constrain the network, further enhancing the performance of pose alignment. In the future, we will try to extend our solution for 3D generation tasks (Di et al. 2024; Luo et al. 2025).
6 Acknowledgments
We acknowledge the support of the advanced computing resources provided by the Supercomputing Center of the USTC.